OpenStudy (anonymous):

. Use a proportion to solve 90 is what percent of 120?

OpenStudy (king):

75 %

OpenStudy (anonymous):

120 - 100% 90 - x% x=100*90/120

OpenStudy (anonymous):

90/120 x 100 =3/4x100=75%